# Compactifications of phylogenetic systems and electrical networks

### v2

@@ 2.2. Phylogenetic Split Systems
A split of $[n]$ is a bipartition $A|B$ of $[n]$ and a split system is any collection of splits of $[n]$. A graph with some of its nodes labeled by $[n]$ displays a split $A|B$ if there is a set of edges whose removal increases the number of connected components by one, and the two new components include respectively the nodes labeled by $A$ and $B$. A minimal display of the split is a displaying set of edges that do not contain a proper subset displaying that split. A split network is a representation of a split system as a graph, where each split is minimally displayed by a set of parallel edges of the same length. A circular split system can be drawn with the $n$ boundary nodes on the boundary of a disk, with non-crossing sets of parallel edges. The following space appears as a chamber in a larger theory of networks and trees in [11]:
@@
Let $\mathbf{\Psi}_{n}$ be the space of weighted circular split systems on $n$ boundary nodes, using the clockwise circular ordering of $[n]$.
The space $\mathbf{\Psi}_{n}$ is stratified with cells corresponding to split systems with no edge weighting. The total number of cells is $|\mathbf{\Psi}_{n}|=2^{n\choose 2}$. A cell $x$ is contained in another cell $y$ when the system corresponding to $x$ can be found by removing splits of $y$. In fact, as a CW-complex $\mathbf{\Psi}_{n}$ is equivalent to a positive orthant of ($\mathbb{R}^{n\choose 2}$. The right side of Figure 3 shows the cells of $\mathbf{\Psi}_{4}$, each column corresponds to cells of a fixed dimension, with $f$-vector $\langle 1,6,15,20,15,6,1\rangle$, the 6th row of Pascal’s triangle.
@@
A circular split system can also be represented by a polygonal diagram [11]: the $n$ terminals label the edges of an $n$-gon in some circular order, and each split is drawn as a diagonal. Figure 2 displays a polygonal diagram in (c) of the circular split system drawing in (d).
To any circular split system $n$, we can associate a dissimilarity matrix, an $n\times n$ real, symmetric, nonnegative matrix, where the $(ij)$-entry is the sum of the weighted splits between nodes $i$ and $j$ in the network. In our case, the dissimilarity matrix is the resistance matrix $W$. It is well-known [36] that there is a bijection between weighted circular split systems and dissimilarity matrices satisfying the Kalmanson condition: there exists a circular order of the boundary such that for any four nodes $i,j,k,l$ listed in that circular order,
@@ 3.1. Graphical Map $\mathbf{\xi}$
We define a key function which takes any equivalence class of circular electrical networks as input and returns a circular planar split system. The input circular networks (and their response matrices) must be with respect to a given circular ordering, typically the clockwise order. While all our maps are defined on weighted networks (edges weighted with conductance and output weighted split systems), we often do abuse notation and apply the maps to unweighted input and output to see their purely combinatorial function.
@@ 3.2. Kalmanson Map $\mathbf{\rho}$
The Kalmanson map $\mathbf{\rho}$ can be extended to all (non-planar) electrical networks by letting even non-Kalmanson resistance metrics be subjected to the best-fit approximation of Neighbor-net. In that extension, the map will no longer be injective. Note also that the algorithm will find a circular order for which $W(N)$ is Kalmanson if possible. This output will only have the original circular order if the original is circular planar in that order.
@@ 3.3. Induced Map $\mathbf{\sigma}$
The weight of a split displayed by a connected circular planar network $N$ is the sum of the weights of 2-groves that respect it, divided by the summed weights of the spanning trees of $N$. The induced split system $\mathbf{\sigma}(N)$ is the set of weighted splits displayed by $N$ [^6] Our $\mathbf{\sigma}(N)$ has the same (unweighted) split network as the map $\Sigma(N)$ from Gambette [16, 22].
@@ 3.4. Matchings and Duals
We construct the dual $N^{*}$ of a planar circular electrical network $N$ using strand matchings. [^7] Another alternative replaces each edge with a perpendicular edge; this method is more difficult to extend to the compactified cactus networks, however. We begin by constructing the medial graph of $N$. Each original edge of $N$ becomes a new node, and any two new nodes arising from adjacent original edges of $N$ are connected by a new edge of the medial graph. Original nodes at the boundary of $N$ are always encircled by one of these new edges; and if the boundary node is degree one then the new edge will be a loop. However, we truncate the medial graph by deleting the portions that extend outside the boundary of $N$, leaving instead a pair of new nodes called stubs on either side of each original boundary node.
The strand diagram of $N$ is found by tracing paths (called strands) in the medial graph, one starting from each stub, and turning neither left nor right at new nodes, but taking the straight option to arrive eventually at another stub. A perfect matching $P$ on $[2n]$ is a set of $n$ pairs $\{a,b\}$ where each element of $[2n]$ is used once. Every circular planar electrical network $N$ gives rise to a perfect matching on $[2n]$, by following its strands. However, this matching depends on $N$ and is not an invariant of the electrical equivalence class. To get an invariant matching, we need to restrict to the critical (reduced) representatives of that class [33].
@@ 3.5. Summary of Maps
Figure 6 displays an overview all the maps discussed so far. We note that maps $\mathbf{\xi}$ and $\mathbf{\xi}^{\prime}$ take as input any electrical network (regardless of planarity) and return a circular planar split system. However, these two maps rely on a given circular order of the boundary nodes, typically given as the clockwise order. In contrast, the map $\mathbf{\sigma}$ must take as input a circular planar electrical network $N$, and yet gives the same result if applied to an equivalent network with a different circular order. Finally, the map $\mathbf{\rho}$ takes as input any electrical network $N$ — Kalmanson, planar, or neither, allowing for imperfect data — and returns a circular planar split system. However this output is only guaranteed to have resistances corresponding to the original network in the case that $N$ has a Kalmanson resistance metric.
@@
Remark.
Noting that a circular planar electrical network can have several consistent circular orders raises the question of a larger space of networks with any circular order allowed. We save the study of such spaces for future work.
@@ 4.1. Cactii
A printed electrical circuit is made of conductors in a network, with terminal nodes on the exterior available for connections. Some of those exterior nodes are identical, short-circuited together so that they are interchangeable in terms of their use. Similarly, monozygotic twins can be seen as genetically the same for the purpose of reconstructing a family tree from the DNA samples of living individuals. On a larger scale, basic DNA sequences of extant individuals from a biologically diverse ecosystem can be equated when the species match. From our perspective, allowing nodes to be infinitely close to each other corresponds to compactifying spaces. Here we define and relate the compactification of phylogenetic split systems (from the DNA examples) to the well-studied compactification of circular electrical network space.
@@
Two cactus networks are equivalent if they have the same response matrix. Lam [33] shows that the (unweighted) electrical equivalence classes of circular planar electrical networks, including those formed by identifying terminal points, correspond bijectively to the perfect matchings on $[2n]$. To find the perfect matching guaranteed by this bijection, the equivalence class must be represented by a reduced $N$. In that case, the strands of the strand diagram will obey the requirement that any two of them cross each other at most once.
@@
For a given circular order on $[n]$, the space of equivalence classes of cactus networks is the compactification ${\mathbf{\Omega}}_{n}$ of the space of circular electrical networks $\mathbf{\Omega}_{n}$. [^9] In [33], the space ${\mathbf{\Omega}}_{n}$ is denoted as $E_{n}$.
@@
Figure 8(a) illustrates the complex of unweighted equivalence classes of cactus networks for $n=3.$ Since the face poset ${\mathbf{\Omega}}_{n}$ is isomorphic to the poset $P_{n}$ of perfect matchings, $|{\mathbf{\Omega}}_{n}|=(2n-1)!!\,.$! .
Figure 8. The graphical map $\mathbf{\xi}$ from cactus networks ${\mathbf{\Omega}}_{3}$ to compactified split systems $\mathbf{\Psi}_{3}$
@@ 4.2. Compactifying Split Systems
Thus, a compactified split system for a given circular order of $[n]$ is a noncrossing partition on $[n]$ and a weighted split system on each part of that partition. Figure 10 shows the associated compactified split system to the network $N^{*}$ from Figure 7, with splits appropriately color-coded.
@@
The space $\mathbf{\Psi}_{n}$ is the set of weighted compactified circular split systems with circular order $[n]$. It is the set of Kalmanson metrics making up the blocks of $n\times n$ dissimilarity matrices $W$.
The unweighted compactified split systems on $[n]$ correspond to the cells of a CW-complex structure on $\mathbf{\Psi}_{n}$. Containment of cells corresponds to either deletion or contraction (via infinity) of splits. The dimension of the top-dimensional cells of $\mathbf{\Psi}_{n}$ is (${n\choose 2}$, since that is the maximum number of splits compatible with a single circular order.
@@
Figure 10. The compactified split system $\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*})$ associated to the network from Figure 7.
Figure 11. The 1-skeleton of $\mathbf{\Psi}_{4}$.
@@ 4.3. Maps
The Kron reduction $K(N)$ of a cactus network can be performed one bulb at a time, or expanded in a disk to show edges with infinite conductance. This expansion allows the edges of $K(N)$ to correspond to the entries in the response matrix $M(N)$. Note that when there is an edge of $N$ incident on a (compactified) vertex (with multiple labels), the edges in $K(N)$ — and thus the entries in $M(N)$ — must evenly divide that conductance. Moreover, the response matrix of a cactus network will have entries of $\infty$ for the conductance between nodes that share a vertex, and corresponding diagonal entries of .$\infty$.- ∞ . The resistance matrix will have entries of 0 for pairs of nodes that share a vertex of the cactus network, but the other entries can be calculated directly by isolating the two nodes and using Kirchoff and Ohm’s laws to find effective resistance between them.
@@
Figure 12 shows a cactus network $N$ and its dual $N^{*}$ together with their Kron reductions and split systems. For instance the edge from 1 to the vertex labeled 2,5,6 has conductance 1; therefore the edges $\{1,2\},\{1,5\},\{1,6\}$ all have conductance $1/3$ in $K(N)$ as seen in the response matrix here:
@@
Figure 12. Cactus networks, duals, and Kron reductions.
@@
First then, the mapping of combinatorial structure can be treated just as for the single-bulb (ordinary) circular planar electrical networks, one bulb of the cactus network at a time. The same is true for electrical networks with more than one connected component: each component can be seen to contribute its own collection of splits on its own boundary nodes. For $N$ a cactus network, the response of the (disconnected) dual network $M(N^{*})$ will decompose into blocks on the diagonal, one for each bulb of $N$, with other entries zero. The resistance matrix $W(N^{*})$ will have corresponding blocks of the same size but with other entries $\infty$.∞ . The maps $\mathbf{\xi}^{\prime}(N^{*})$ and $\mathbf{\rho}(N^{*})$ can thus operate block-wise.
To see that this block-wise operation gives the expected values of the weights of the splits, we note that the process of compactifying a split system (via treating some splits as infinite) sums the weights of the splits which coincide after collapsing infinite edges to a point; see Figure 9. Thus the weight of the final split in the compactified system $\mathbf{\xi}(N)$ is the same as the edge in the Kron reduction of a single bulb, which is the weight of the split found by $\mathbf{\rho}(N^{*})$ as in the connected case (Theorem 13). ∎
@@ 5.1. Lagrange Inversion
Our graphical map $\mathbf{\xi}$ takes any compactified circular network to a compactified planar split system. We leverage this map and show how to count these systems, all based on the following result:
@@
For a given circular order on $[n]$, the number of unweighted compactified split systems is given by the OEIS sequence A136654. The generating function is given by
@@
To find the number of compactified split systems on $n,$ use the Lagrange inversion theorem on $x/B(x)$: Find the series expansion for $(B_{n}(x))^{n+1}$ (where $B_{n}(x)$ is the partial series up to $x^{n}$), take the coefficient of $x^{n}$ in that expansion, and divide by $n+1$. For instance, when $n=4$, we see
@@
We start by associating a species $A$ to the combinatorial structures of compactified split systems. Define $A([n])$ to be the set of pairs made up of a compactified split system on $[n]$ and a permutation on $[n].$ The permutation can be thought of as a 1-1 mapping of the exterior labels in their cyclic order to a new set of numerical tags $\{1,\dots,n\}$. This allows us to consider $A(x)$ as simultaneously the ordinary generating function of the numbers of structure (types) of compactified split systems and the exponential generating function of the species $A.$
Next, any instance of this species on $[n]$ can be created with the following steps. First we choose $k$ tags: any subset of $[n]$ of size $k$, and also a subset $S_{1}$ of our original cyclic labels of size $k$ but required to contain the label 1.111111For instance, to make the structure in Figure 10 we let $k=2$ and chose $S_{1}=\{1,15\}.$ The tags are not shown in Figure 10 because we only wanted to display the structure type. The tags are attached to the labels in $S_{1}$ (in any order) and a set of splits is assigned to $S_{1}.$ There are $2^{k\choose 2}$ possibilities for those splits.
Set $S_{1}$ is a part of a noncrossing partition of the circular labels $[n]$, and the remaining labels not in $S_{1}$ are subdivided by the contiguous portions of $S_{1}$ into at most $k$ contiguous (in the circular order) sets. These sets are then used to continue the process: the species $A$ is applied to each of them recursively, with all the possible partitions of the remaining $n-k$ tags. Thus we see a functional equation for the species:
@@
or equivalently, $A=B\circ(X\cdot A).$ Then in terms of exponential generating functions, $A(x)=B(xA(x)).$ Multiplying both sides by $x$ and rearranging yields

### v3

@@ 2.2. Phylogenetic Split Systems
A split of $[n]$ is a bipartition $A|B$ of $[n]$ and a split system is any collection of splits of $[n]$. A graph with some of its nodes labeled by $[n]$ displays a split $A|B$ if there is a set of edges whose removal increases the number of connected components by one, and the two new components include respectively the nodes labeled by $A$ and $B$. A minimal display of the split is a displaying set of edges that do not contain a proper subset displaying that split. A split network is a representation of a split system as a graph, where each split is minimally displayed by a set of parallel edges of the same length. A circular split system can be drawn as a split network with the $n$ boundary nodes on the boundary of a disk, and with non-crossing sets of parallel edges. The following space appears as a single chamber in a larger space of networks, those allowing any ordering of their boundaries, in [11]:
@@
Let $\mathbf{\Psi}_{n}$ be the space of weighted circular split systems on $n$ boundary nodes, using the clockwise cyclic ordering of $[n]$.
The space $\mathbf{\Psi}_{n}$ is stratified with cells corresponding to split systems with no edge weighting. The total number of cells is $|\mathbf{\Psi}_{n}|=2^{n\choose 2}$. A cell $x$ is contained in another cell $y$ when the system corresponding to $x$ can be found by removing splits of $y$. In fact, as a CW-complex $\mathbf{\Psi}_{n}$ is equivalent to a positive orthant of ($\mathbb{R}^{n\choose 2}$. The right side of Figure 4 shows the cells of $\mathbf{\Psi}_{4}$, each column corresponds to cells of a fixed dimension, with $f$-vector $(1,6,15,20,15,6,1)$, the 6th row of Pascal’s triangle.
@@
A circular split system can also be represented by a polygonal diagram [11]: the $n$ terminals label the edges of an $n$-gon in some cyclic order, and each split is drawn as a diagonal. Figure 3 displays a polygonal diagram in (c) of the circular split network drawing in (d).
To any circular split system $n$, we can associate a dissimilarity matrix, an $n\times n$ real, symmetric, nonnegative matrix, where the $(ij)$-entry is the sum of the weighted splits between nodes $i$ and $j$ in the network. In our case, the dissimilarity matrix is the resistance matrix $W$. It is well-known [38] that there is a bijection between weighted circular split systems and dissimilarity matrices satisfying the Kalmanson condition: there exists a cyclic order of the boundary such that for any four nodes $i,j,k,l$ listed in that cyclic order,
@@ 3.1. Graphical Map $\mathbf{\xi}$
We define a key function which takes any equivalence class of circular electrical networks as input and returns a circular planar split system. The input circular networks (and their response matrices) must be with respect to a given cyclic ordering, typically the clockwise order. While all our maps are defined on weighted networks (edges weighted with conductance and output weighted split systems), we often do abuse notation and apply the maps to unweighted input and output to see their purely combinatorial function.
@@ 3.2. Kalmanson Map $\mathbf{\rho}$
The Kalmanson map $\mathbf{\rho}$ can be extended to all (non-planar) electrical networks by letting even non-Kalmanson resistance metrics be subjected to the best-fit approximation of Neighbor-net. In that extension, the map will no longer be injective. Note also that the algorithm will find a cyclic order for which $W(N)$ is Kalmanson if possible. This output will only have the original cyclic order if the original is circular planar in that order.
@@ 3.3. Induced Map $\mathbf{\sigma}$
The weight of a split displayed by a connected circular planar network $N$ is the sum of the weights of 2-groves that respect it, divided by the summed weights of the spanning trees of $N$. The induced split system $\mathbf{\sigma}(N)$ is the set of weighted splits displayed by $N$ [^5] Our $\mathbf{\sigma}(N)$ has the same underlying unweighted split network as given by the map $\Sigma(N)$ from Gambette [15, 21].
@@ 3.4. Matchings and Duals
We construct the dual $N^{*}$ of a planar circular electrical network $N$ using strand matchings. [^6] Another alternative replaces each edge with a perpendicular edge; however this method is more difficult to extend to the cactus networks. We begin by constructing the medial graph of $N$. Each original edge of $N$ becomes a new node, and any two new nodes arising from adjacent original edges of $N$ are connected by a new edge of the medial graph. Original nodes at the boundary of $N$ are always encircled by one of these new edges; and if the boundary node is degree one then the new edge will be a loop. However, we truncate the medial graph by deleting the portions that extend outside the boundary of $N$, leaving instead a pair of new nodes called stubs on either side of each original boundary node.
The strand diagram of $N$ is found by tracing paths (called strands) in the medial graph, one starting from each stub, and turning neither left nor right at new nodes, but taking the straight option to arrive eventually at another stub. A perfect matching $P$ on $[2n]$ is a set of $n$ pairs $\{a,b\}$ where each element of $[2n]$ is used once. Every circular planar electrical network $N$ gives rise to a perfect matching on $[2n]$, by following its strands. However, this matching depends on $N$ and is not an invariant of the electrical equivalence class. To get an invariant matching, we need to restrict to the critical (reduced) representatives of that class as in [34], which can be recognized from the fact that strands never cross each other more than once (lens-free diagrams). In that source Lam shows that the critical representatives are in bijection with all $(2n-1)!!$! perfect matchings. Here we use the strands of even non-critical networks (as in our running example from Figure 2) to find the dual, and point out that Kron reductions of $N$ and its dual are invariants of the electrical equivalence classes. (So our maps’ dependence only on the Kron reduction proves, for instance, that the calculation of the weight of a split will yield the same result no matter what specific $N$ represents the equivalence class.)
@@ 3.5. Summary of Maps
Figure 7 displays an overview all the maps discussed so far. We note that maps $\mathbf{\xi}$ and $\mathbf{\xi}^{\prime}$ take as input any electrical network (regardless of planarity) and return a circular planar split system. However, these two maps rely on a given cyclic order of the boundary nodes, typically given as the clockwise order. In contrast, the map $\mathbf{\sigma}$ must take as input a circular planar electrical network $N$, and yet gives the same result if applied to an equivalent network with a different cyclic order. Finally, the map $\mathbf{\rho}$ takes as input any electrical network $N$ — Kalmanson, planar, or neither, allowing for imperfect data — and returns a circular planar split system. However this output is only guaranteed to have resistances corresponding to the original network in the case that $N$ has a Kalmanson resistance metric.
@@ 4.1. Cactii
Two cactus networks are equivalent if they have the same sets of response matrices, one for each bulb. Lam [34] shows that the (unweighted) electrical equivalence classes of circular planar electrical networks, including those formed by identifying terminal points, correspond bijectively to the perfect matchings on $[2n]$. To find the perfect matching guaranteed by this bijection, the equivalence class must be represented by a reduced $N$. In that case, the strands of the strand diagram will obey the requirement that any two of them cross each other at most once.
@@
For a given cyclic order on $[n]$, the space of equivalence classes of cactus networks is the compactification ${\mathbf{\Omega}}_{n}$ of the space of circular electrical networks $\mathbf{\Omega}_{n}$. [^8] In [34], the space ${\mathbf{\Omega}}_{n}$ is denoted as $E_{n}$.
@@
Figure 9(a) illustrates the complex of unweighted equivalence classes of cactus networks for $n=3,$ with $f$-vector (5, 6, 3, 1). Figure 13 shows the 1-skeleton of the complex ${\mathbf{\Omega}}_{4}$ (with cells labeled by split systems). The $f$-vector for the complex ${\mathbf{\Omega}}_{4}$ is (14, 28, 28, 20, 10, 4, 4, 1). Since the face poset ${\mathbf{\Omega}}_{n}$ is isomorphic to the poset $P_{n}$ of perfect matchings, $|{\mathbf{\Omega}}_{n}|=(2n-1)!!\,.$! .
Figure 9. The graphical map $\mathbf{\xi}$ from cactus electrical networks ${\mathbf{\Omega}}_{3}$ to cactus split systems $\mathbf{\Psi}_{3}$
@@ 4.2. Compactifying Split Systems
Thus, a cactus split system for a given cyclic order (clockwise) of $[n]$ is a non-crossing partition on $[n]$ and a weighted split system on each part of that partition. Figure 12 shows the associated cactus split system to the network $N^{*}$ from Figure 8, with splits appropriately color-coded.
@@
The space $\mathbf{\Psi}_{n}$ is the set of weighted cactus circular split systems with clockwise cyclic order of $[n]$. Thus it is the set of Kalmanson metrics making up the blocks of $n\times n$ dissimilarity matrices $W$, where blocks correspond to the parts of the non-crossing partition of $[n]$.
The unweighted cactus split systems on $[n]$ correspond to the cells of a CW-complex structure on $\mathbf{\Psi}_{n}$. Containment of cells corresponds to either deletion or contraction (via infinity) of splits. The dimension of the top-dimensional cells of $\mathbf{\Psi}_{n}$ is (${n\choose 2}$, since that is the maximum number of splits compatible with a single cyclic order.
@@ 4.3. Maps
The Kron reduction $K(N)$ of a cactus network can be performed one bulb at a time. This is the same procedure as considering the response matrices of the decomposition of $N$ into one network for each bulb, embedded in its own disk, as in the proof of Theorem 4.9 in [33].
Remark.
In [33] is is stated that “There is a natural notion of the response matrix of a cactus network: we specify voltages at boundary vertices such that identified vertices are assigned the same voltage.” This matrix $M(N)$ can be posited as the limit of the response matrix as some of the edges of a network are taken to zero resistance. Also the Kron reduction of $N$ can be seen as expanded in a single disk to show edges with infinite conductance. Note that when there is an edge of $N$ incident on a (compactified) vertex (with multiple labels), the edges in $K(N)$ — and thus the entries in $M(N)$ — must evenly divide that conductance. Moreover, the response matrix of a cactus network will have entries of $\infty$ for the conductance between nodes that share a vertex, and corresponding diagonal entries of .$\infty$.- ∞ . The corresponding resistance matrix will have entries of 0 for pairs of nodes that share a vertex of the cactus network, but the other entries can be calculated directly by isolating the two nodes and using Kirchoff and Ohm’s laws to find effective resistance between them.
@@
Figure 11 shows a cactus network $N$ and its dual $N^{*}$ together with their associated split systems. Each bulb of $N$ is its own Kron reduction. Also we can find a limiting $M(N)$ and $W(N)$ for the entire network, while here $M(N^{*})$ and $W(N^{*})$ exhibit block matrix structure.
@@
Figure 11. Cactus networks, duals, Kron reductions and corresponding split systems.
@@
First then, the mapping of combinatorial structure can be treated just as for the single-bulb (ordinary) circular planar electrical networks, one bulb of the cactus network at a time. The same is true for electrical networks with more than one connected component: each component can be seen to contribute its own collection of splits on its own boundary nodes. For $N$ (or $N^{*}$) a cactus network, we can construct small response matrices separately for each bulb of $N$, choosing a name for the identified nodes on the boundary of that bulb. Then each of those bulbs will also have a resistance matrix of the same size. The maps $\mathbf{\xi}^{\prime}(N^{*})$ and $\mathbf{\rho}(N^{*})$ can thus operate bulb-wise. To see that this operation gives the expected values of the weights of the splits, we note that the process of compactifying a split system (via treating some splits as infinite) sums the weights of the splits which coincide after collapsing any edges to a point; see Figure 10. Thus the weight of the final split in the cactus split system $\mathbf{\xi}(N)$ is the same as the edge in the Kron reduction of a single bulb, which is the weight of the split found by $\mathbf{\rho}(N^{*})$ as in the connected case (Theorem 13). ∎
Figure 12. The cactus split system $\mathbf{\sigma}(N)=\mathbf{\xi}^{\prime}(N^{*})$ associated to the network from Figure 8.
Figure 13. The 1-skeleton of $\mathbf{\Omega}_{4}$. Each 0-cell has 3,4, or 6 attached labeled 1-cells.
Figure 14. The 1-skeleton of $\mathbf{\Psi}_{4}$. Cells are images of $\mathbf{\xi}^{\prime}$ from Figure 13.
@@ 5.1. Lagrange Inversion
Our graphical map $\mathbf{\xi}$ takes any cactus circular network to a cactus planar split system. We leverage this map and show how to count these systems, all based on the following result:
@@
For a given cyclic order on $[n]$, the number of unweighted cactus split systems is given by the OEIS sequence A136654. The generating function is given by
@@
To find the number of cactus split systems on $n,$ use the Lagrange inversion theorem on $x/B(x)$: Find the series expansion for $(B_{n}(x))^{n+1}$ (where $B_{n}(x)$ is the partial series up to $x^{n}$), take the coefficient of $x^{n}$ in that expansion, and divide by $n+1$. For instance, when $n=4$, we see
@@
We start by associating a species $A$ to the combinatorial structures of cactus split systems. Define $A([n])$ to be the set of pairs made up of a cactus split system on $[n]$ and a permutation on $[n].$ The permutation can be thought of as a 1-1 mapping of the exterior labels in their cyclic order to a new set of numerical tags $\{1,\dots,n\}$. This allows us to consider $A(x)$ as simultaneously the ordinary generating function of the numbers of structure (types) of cactus split systems and the exponential generating function of the species $A.$
Next, any instance of this species on $[n]$ can be created with the following steps. First we choose $k$ tags: any subset of $[n]$ of size $k$, and also a subset $S_{1}$ of our original cyclic labels of size $k$ but required to contain the label 1.101010For instance, to make the structure in Figure 12 we let $k=2$ and chose $S_{1}=\{1,15\}.$ The tags are not shown in Figure 12 because we only wanted to display the structure type. The tags are attached to the labels in $S_{1}$ (in any order) and a set of splits is assigned to $S_{1}.$ There are $2^{k\choose 2}$ possibilities for those splits.
Set $S_{1}$ is a part of a noncrossing partition of the circular labels $[n]$, and the remaining labels not in $S_{1}$ are subdivided by the contiguous portions of $S_{1}$ into at most $k$ contiguous (in the cyclic order) sets. These sets are then used to continue the process: the species $A$ is applied to each of them recursively, with all the possible partitions of the remaining $n-k$ tags. Thus we see a functional equation for the species:
@@
or equivalently, $A=B\circ(X\cdot A)$, where $B$ is the species of ordinary split systems and $X$ is the singleton species. Then in terms of exponential generating functions, $A(x)=B(xA(x))$ Multiplying both sides by $x$ and rearranging yields